OpenStudy (anonymous):

multiply/simplify 25x^2/4x^2-32x+64 * 4x-16/5x

OpenStudy (saifoo.khan):

\[\frac{5x}{x-4}\]

OpenStudy (anonymous):

\[\frac{25x^2}{4x^2-32x+64}\cdot\frac{4x-16}{5x}=\] \[\frac{25x^2}{4(x-4)^2}\cdot\frac{4(x-4)}{5x}=\] \[\frac{5x}{x-4}\]

OpenStudy (saifoo.khan):

i won! =D